Should NOT print DRAFT since, conditional is used.

\usepackage[conditional]{draftcopy}

Test 13 This is a test page 1

Test 13 This is a test page 2

This is a testpage 1 This is a testpage 2 This is a testpage 3 This is a testpage 4